# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ B.3 End of the proof of the induction step
with $D_{0}=\frac{\partial}{\partial t}$ and $D_{m}$ as in (61) for $m>0$. We plug (152) into the above relation. Since $D_{\ell_{n+1}}$ is a differential operator that annihilates the variable $t$ except when $\ell_{n+1}=0$, the first term is equal to
$D_{\ell_{n+1}}\left(R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\right)=\frac{\ell_{1}+\cdots+\ell_{n}}{2}\cdot R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\frac{D_{\ell_{n+1}}R}{R}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\\ +R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\sum_{r=1}^{d_{g,n}}\left(D_{\ell_{n+1}}\left(\frac{R^{(r)}}{R}\right)\frac{\partial P^{(g)}_{n,I}}{\partial x_{r}}(\ell_{1}^{2},\ldots,\ell_{n}^{2})+D_{\ell_{n+1}}\left(\frac{S^{(r)}}{\sqrt{R}}\right)\frac{\partial P^{(g)}_{n,I}}{\partial y_{r}}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\right)\\ -c(g,n+1)t^{2-2g-(n+1)}\delta_{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1},0}$ (154)
where we set $c(g,n+1)=(2-2g-n)c(g,n)$. By Proposition B.3 and its consequence discussed around (149), the second term of (154) is of the wanted form $R^{\frac{\ell_{1}+\cdots+\ell_{n+1}}{2}}$ times a quasi-polynomial in $\ell_{1}^{2},\ldots,\ell_{n+1}^{2}$. We then split the second term of (153) into pieces.
Assume first that that $\ell_{n+1}$ is even, in which case, the quantity $D_{\ell_{n+1}}R/R$ appearing in the first term of (154) equals $R^{\ell_{n+1}/2}R^{\prime}/R$. The last term of (153) is then
@@ Appendix C Quasi-polynomiality in the general case: initialisation
That the statement holds in the planar case $(g,n)=(0,3)$ is a direct consequence of Theorem A.1 and of the relation (23) between the $T$’s and the $\tau$’s. To treat the case of higher genera, we will need some results from the theory of topological recursion presented in [Eynard2016], and more precisely, we will need the expression of the generating function of maps in genus $g\geq 1$ based on the so-called method of moments [Ambjoern1995]. Our approach can be seen as an extension of the analysis of Budd [budd2020irreducible, Section 3.4] to the non bipartite case (but only considering the non irreducible case). Recall that $F^{(g)}=F^{(g)}_{\varnothing}=T^{(g)}_{\varnothing}=\tau^{(g)}_{\varnothing}$ denote the generating series of maps of genus $g$ without boundaries.
@@ C.1 Genus one
Let us recall Eynard’s notation from [Eynard2016] (Eynard puts a bound on the degrees, and excludes faces of degrees $1$ and $2$, we are going to ignore this). We set (p.58 and Theorem 3.1.2 p.62) $x(z)=S+\sqrt{R}(z+1/z)$ ($S=\alpha,R=\gamma^{2}$), and (pp.59-60, Definition 3.1.1 p.63 and Theorem 3.1.2 p.62)
@@
$u_{k}=S\delta_{k,0}+\sqrt{R}\delta_{k,1}-\sum_{i\geq k+1}t_{i}\sum_{j=k}^{[(i+k)/2]-1}\binom{i-1}{j,j-k,i-1+k-2j}R^{j-k/2}S^{i-1+k-2j}\,.$ (167)
Note that the latter is related to the generating function $V_{k-1}$ for $(k-1)$-slices introduced in [irredmaps]: for $k\geq 2$, we have $V_{k-1}=-R^{k/2}u_{k}$. On the other hand, we have $u_{0}=0$ and $u_{1}=t/\sqrt{R}$, which amounts to (3). We define the 00-th order moments by
$M_{+,0}=\frac{-y^{\prime}(1)}{\sqrt{R}}\,,\qquad M_{+,0}=\frac{-y^{\prime}(-1)}{\sqrt{R}}.$ (168)
Now, [Eynard2016, Theorem 3.4.6] expresses the generating function of toric maps in terms of $M_{\pm,0}$ as
@@
On the other hand, [Eynard2016, Theorem 3.3.4] gives
$\begin{split}\omega_{3}^{(0)}(z_{1},z_{2},z_{3})=&\frac{1}{2RM_{+,0}}\frac{1}{(z_{1}-1)^{2}(z_{2}-1)^{2}(z_{3}-1)^{2}}\\ &-\frac{1}{2RM_{-,0}}\frac{1}{(z_{1}+1)^{2}(z_{2}+1)^{2}(z_{3}+1)^{2}}\,,\end{split}$ (170)
@@ C.2 Higher moments, higher genera
The generating functions for maps in higher genera are given by higher order moments, which are given ([Eynard2016, p.64]), for $h\geq 1$, by
@@
One should note that [Eynard2016, Corollary 3.5.1] introduces the quantities $R^{h/2}M_{\pm,h}$ that are called “dimensionless”, we will see that these are indeed natural quantities to consider, due to the following result [^8] There seems to be an inaccuracy in the statement of [Eynard2016, Corollary 3.5.1]. Indeed, the latter is stated without the boundary term involving $t^{2-2g}$ only, which is however present in other similar statements such as [ibid., Theorems 3.4.3 and 3.4.9]. In these statements, the coefficient $\frac{B_{2g}}{2g(2-2g)}$ is indeed equal to $\chi(\mathcal{M}_{g,0})$ as wanted. It seems that this boundary term got omitted at some stage.:
@@
[Eynard2016, Corollary 3.5.1] There exists $\mathcal{P}_{g}\in\mathbb{Q}[x_{+},x_{-},(y_{+,h},y_{-,h},1\leq h\leq 3g-3)]$ that is homogeneous of degree $22g-$ 2 in the first two variables, and such that $\mathcal{P}_{g}(x_{+},x_{-},(\mu^{h},\mu^{h},1\leq h\leq 3g-3))$ is of degree $33g-$ 3 in $\mu$, and $c(g,0)\in\mathbb{Q}$, such that
@@
We would like to show that—as for $M_{\pm,0}$, recall (173)—the “dimensionless” moments $R^{h/2}M_{\pm,h}$ are rational functions of $R^{(k)}/R,S^{(k)}/\sqrt{R}$ for $11\leq k\leq h+$ 1. Our goal will be to get rid of the variables $t_{i}$ appearing implicitly in the $u_{k}$ in the previous expression, and express this purely in terms of $R,S$ and their derivatives. Our approach is inspired by that of [budd2020irreducible, Section 3.4].
@@
$\bar{M}_{+,h}=\sum_{i\geq 2}t_{i}\sum_{k}\binom{k+h}{2h+1}\sum_{j}\binom{i-1}{j,j-k,i-1+k-2j}R^{j-k/2}S^{i-1+k-2j}\\ =\sum_{i\geq 2}t_{i}\sum_{l}R^{\frac{i-1-l}{2}}S^{l}\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}$ (177)
where the second expression comes from substituting the summation over $j$ to the summation over $l=i-1+k-2j$. To alleviate notations, we do not specify the summation ranges for $k,j,l$ as they are naturally enforced by the vanishing of the binomial and multinomial coefficients and by the requirement that their arguments are nonnegative integers. Note that, in the summation over $i$, the terms $i=2,\ldots,h+1$ have a vanishing contribution.
@@ C.2.1 A binomial identity
We now state the following variant of [budd2020irreducible, Lemma 7].
@@
For any nonnegative integer $h$, there exist two polynomials $Q_{h}^{[0]}$ and $Q_{h}^{[1]}$ of degree $h+1$ such that, denoting by $\epsilon=\epsilon(i,l)\in\{0,1\}$ the parity of $i+l$, we have
$\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}=\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q_{h}^{[\epsilon]}\left(\frac{i-l-\epsilon}{2}\right)$ (178)
@@
Proof.
We rewrite the sum as
$\binom{i-1}{l}\sum_{k}\binom{k+h}{2h+1}\binom{i-1-l}{\frac{i-1+k-l}{2}}$ (179)
Assume that $i-l$ is even, that is, $\epsilon=0$, so that the summation is over odd $k=2k^{\prime}+1$. Then, writing $2j^{\prime}=i-l$, this rewrites
$\binom{i-1}{l}\sum_{k^{\prime}}\binom{2k^{\prime}+h+1}{2h+1}\binom{2j^{\prime}-1}{j^{\prime}+k^{\prime}}=\binom{i-1}{l}\binom{2j^{\prime}-1}{j^{\prime}}Q_{h}(j^{\prime})\,,$ (180)
where $Q_{h}=Q_{h}(0,\cdot)$ is the polynomial of degree $h+1$ that appears in [budd2020irreducible, Lemma 7]. Alternatively, we can derive this lemma by the following independent argument. First note that, in the notation of [polytightmaps], we have
@@
$\mathsf{A}_{j,k}=\frac{2k}{2j}\binom{2j}{j+k}\,,\quad j,k\in\frac{1}{2}\mathbb{Z}$ (182)
from [polytightmaps, Equation (3.2)], we can rewrite the left hand side of (180) in the form
$\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}2(k^{\prime}+1/2)p_{h}(2(k^{\prime}+1/2))\binom{2(j^{\prime}-1/2)}{j^{\prime}-1/2+k^{\prime}+1/2}\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}p_{h}(2(k^{\prime}+1/2))\,.$ (183)
@@
The polynomial $p_{h}(2m)$ is of degree $h$ in $m^{2}$, where $m\in\mathbb{Z}+1/2$, so it can be expressed in the basis of the polynomials $\tilde{p}_{r}(m),0\leq r\leq h$ defined in [polytightmaps, Equation (2.18)], in the form $p_{h}(2m)=\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\tilde{p}_{r}(m)$, which, based on [polytightmaps, Equation (3.9)], yields the expression
$(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}\,\tilde{p}_{r}(k^{\prime}+1/2)\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}\,\tilde{p}_{r}(k^{\prime}+1/2)\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\binom{j^{\prime}-1}{r}\binom{2j^{\prime}-2}{j^{\prime}-1}\\ =\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}\frac{h!^{2}}{(2h+1)!}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\,j^{\prime}\binom{j^{\prime}-1}{r}$ (184)
@@
which, since $j^{\prime}=(i-l)/2$, is indeed of the wanted form
$\binom{i-1}{\frac{i-l}{2},\frac{i-l}{2}-1,l}Q_{h}^{[0]}\left(\frac{i-l}{2}\right)\,,$ (185)
for some polynomial $Q^{[0]}_{h}$ of degree $h+1$.
@@
For $\epsilon=1$, we argue in a similar (and slightly simpler) way. The summation is now over even $k=2k^{\prime}$, and we write $12j^{\prime}=i-l-$ 1, so that (179) rewrites, from [polytightmaps, Equation (3.2)]:
$\binom{i-1}{l}\sum_{k^{\prime}}\binom{2k^{\prime}+h}{2h+1}\binom{2j^{\prime}}{j^{\prime}+k^{\prime}}=2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime},k^{\prime}}p_{h}(2k^{\prime})\,.$ (186)
@@
This time, we express the polynomial $p_{h}(2m)$ in the basis of the polynomials $p_{r}(m),0\leq r\leq h$, in the form $p_{h}(2m)=\sum_{r=0}^{h}\alpha_{h,r}p_{r}(k)$, which yields, now using [polytightmaps, Equation (3.7)], the expression
$2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{k^{\prime}}\sum_{r=0}^{h}\alpha_{h,r}\mathsf{A}_{j^{\prime},k^{\prime}}p_{r}(k^{\prime})\\ =2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\alpha_{h,r}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime},k^{\prime}}p_{r}(k^{\prime})\\ =2j^{\prime}\frac{h!^{2}}{(2h+1)!}\binom{i-1}{l}\sum_{r=0}^{h}\alpha_{h,r}\binom{j^{\prime}-1}{r}\binom{2j^{\prime}-1}{j^{\prime}}\\ =\binom{i-1}{\frac{i-l-1}{2},\frac{i-l-1}{2},l}\frac{h!^{2}}{(2h+1)!}j^{\prime}\sum_{r=0}^{h}\alpha_{h,r}\binom{j^{\prime}-1}{r}\,,$ (187)
@@
which is again of the wanted form. ∎
C.2.2 Expressing the moments in terms of derivatives of $\mathbf{Z}$
Lemma C.3 and (177) allow to rewrite the moment $M_{+,h}$ as
$\bar{M}_{+,h}=\sum_{\epsilon\in\{0,1\}}\sum_{i\geq 2}t_{i}\sum_{l}R^{\frac{i-1-l}{2}}S^{l}\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q^{[\epsilon]}_{h}\left(\frac{i-l-\epsilon}{2}\right)\ ,$ (188)
and
$(-1)^{h}\bar{M}_{-,h}=\sum_{\epsilon\in\{0,1\}}(-1)^{\epsilon}\sum_{i\geq 2}t_{i}\sum_{l}R^{\frac{i-1-l}{2}}S^{l}\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q^{[\epsilon]}_{h}\left(\frac{i-l-\epsilon}{2}\right)\ ,$ (189)

### v2

@@ B.3 End of the proof of the induction step
with $D_{0}=\frac{\partial}{\partial t}$ and $D_{m}$ as in (62) for $m>0$. We plug (155) into the above relation. Since $D_{\ell_{n+1}}$ is a differential operator that annihilates the variable $t$ except when $\ell_{n+1}=0$, the first term on the right-hand side of (156) is equal to
$D_{\ell_{n+1}}\left(R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,\ell_{n}^{2})-c(g,n)t^{2-2g-n}\delta_{\ell_{1}+\cdots+\ell_{n},0}\right)=\\ \frac{\ell_{1}+\cdots+\ell_{n}}{2}\cdot R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\frac{D_{\ell_{n+1}}R}{R}P^{(g)}_{n,I}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\\ +R^{\frac{\ell_{1}+\cdots+\ell_{n}}{2}}\sum_{r=1}^{d_{g,n}}\left(D_{\ell_{n+1}}\left(\frac{R^{(r)}}{R}\right)\frac{\partial P^{(g)}_{n,I}}{\partial x_{r}}(\ell_{1}^{2},\ldots,\ell_{n}^{2})+D_{\ell_{n+1}}\left(\frac{S^{(r)}}{\sqrt{R}}\right)\frac{\partial P^{(g)}_{n,I}}{\partial y_{r}}(\ell_{1}^{2},\ldots,\ell_{n}^{2})\right)\\ -c(g,n+1)t^{2-2g-(n+1)}\delta_{\ell_{1}+\cdots+\ell_{n}+\ell_{n+1},0}$ (157)
where we set c(g,n+1)=(22gn)c(g,n)c(g,n+1)=(2-2g-n)c(g,n). By Proposition B.3 and its consequence discussed around (152), the third line of (157) is of the wanted form $R^{\frac{\ell_{1}+\cdots+\ell_{n+1}}{2}}$ times a quasi-polynomial in $\ell_{1}^{2},\ldots,\ell_{n+1}^{2}$. We then split the last term of (156) into pieces.
Assume first that that $\ell_{n+1}$ is even, in which case, the quantity $D_{\ell_{n+1}}R/R$ appearing in the second line of (157) equals $R^{\ell_{n+1}/2}R^{\prime}/R$. The last term of (156) is then
@@ Appendix C Quasi-polynomiality in the general case: initialisation
The fact that the statement holds in the planar case $(g,n)=(0,3)$ is a direct consequence of Theorem A.1 and of the relation (23) between the $T$’s and the $\tau$’s. To treat the case of higher genera, we will need some results from the theory of topological recursion presented in [Eyn16], and more precisely, we will need the expression of the generating function of maps in genus $g\geq 1$ based on the so-called method of moments [ACKM93]. Our approach can be seen as an extension of the analysis of Budd [Bud22, Section 3.4] to the non bipartite case (but only considering the non-irreducible case). Recall that $F^{(g)}=F^{(g)}_{\varnothing}=T^{(g)}_{\varnothing}=\tau^{(g)}_{\varnothing}$ denote the generating series of maps of genus $g$ without boundaries.
@@ C.1 Genus one
Let us recall Eynard’s notation [^9] Eynard puts a bound on the degrees, and excludes faces of degrees $1$ and $2$, but this restriction is in fact inessential, as it results from Eynard’s choice of considering the map generating functions as formal power series in the vertex weight $t$ whose coefficients are polynomials in the face weights. Working instead with a weight per edge lifts this restriction since the set of maps with a given number of edges is finite, and we believe that the reasonings of [Eyn16] extend to this setting. from [Eyn16]. We set (p.58 and Theorem 3.1.2 p.62) $x(z)=S+\sqrt{R}(z+1/z)$ ($S=\alpha,R=\gamma^{2}$), and (pp.59-60, Definition 3.1.1 p.63 and Theorem 3.1.2 p.62)
@@
$u_{k}=S\delta_{k,0}+\sqrt{R}\delta_{k,1}-\sum_{i\geq k+1}t_{i}\sum_{j=k}^{\lfloor(i+k-1)/2\rfloor}\binom{i-1}{j,j-k,i-1+k-2j}R^{j-k/2}S^{i-1+k-2j}\,.$ (170)
Note that the latter is related to the generating function $V_{k-1}$ for (k1)(k-1)-slices introduced in [BG14]: for $k\geq 2$, we have $V_{k-1}=-R^{k/2}u_{k}$. On the other hand, we have $u_{0}=0$ and $u_{1}=t/\sqrt{R}$, as seen by applying the multinomial formulas in (3). We define the 0-th order moments by
$M_{+,0}=\frac{-y^{\prime}(1)}{\sqrt{R}}\,,\qquad M_{-,0}=\frac{-y^{\prime}(-1)}{\sqrt{R}}.$ (171)
Now, [Eyn16, Theorem 3.4.6] expresses the generating function of toric maps in terms of $M_{\pm,0}$ as
@@
On the other hand, [Eyn16, Theorem 3.3.4] gives
[…]
@@ C.2 Higher moments, higher genera
The generating functions for maps in higher genera are given by higher order moments, which are given ([Eyn16, p.64]), for $h\geq 1$, by
@@
One should note that [Eyn16, Corollary 3.5.1] introduces the quantities $R^{h/2}M_{\pm,h}$ that are called “dimensionless”, we will see that these are indeed natural quantities to consider, due to the following result [^10] There seems to be an inaccuracy in the statement of [Eyn16, Corollary 3.5.1]. Indeed, the latter is stated without the boundary term involving $t^{2-2g}$ only, which is however present in other similar statements such as [ibid., Theorems 3.4.3 and 3.4.9]. In these statements, the coefficient $\frac{B_{2g}}{2g(2-2g)}$ is indeed equal to $\chi(\mathcal{M}_{g,0})$ as wanted. It seems that this boundary term got omitted at some stage.:
@@
[Eyn16, Corollary 3.5.1] There exists $\mathcal{P}_{g}\in\mathbb{Q}[x_{+},x_{-},(y_{+,h},y_{-,h},1\leq h\leq 3g-3)]$ that is homogeneous of degree 2g22g-2 in the first two variables, and such that $\mathcal{P}_{g}(x_{+},x_{-},(\mu^{h},\mu^{h},1\leq h\leq 3g-3))$ is of degree 3g33g-3 in $\mu$, and $c(g,0)\in\mathbb{Q}$, such that
@@
We would like to show that, for $h\geq 1$, the “dimensionless” moments $R^{h/2}M_{\pm,h}$ are rational functions of $R^{(k)}/R,S^{(k)}/\sqrt{R}$ for $1\leq k\leq h+1$. Our goal will be to get rid of the variables $t_{i}$ appearing implicitly in the $u_{k}$ in the previous expression, and express this purely in terms of $R,S$ and their derivatives. Our approach is inspired by that of [Bud22, Section 3.4].
@@
$\bar{M}_{\pm,h}=\!\!\sum_{k\geq h+1}\!\!(\pm 1)^{k+h+1}\sum_{i\geq 1}t_{i}\binom{k+h}{2h+1}\sum_{j}\binom{i-1}{j,j-k,i-1+k-2j}R^{j-k/2}S^{i-1+k-2j}\\ =(\pm 1)^{h}\sum_{i\geq h+2}t_{i}\sum_{l}(\pm 1)^{i-l}R^{\frac{i-l-1}{2}}S^{l}\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}$ (180)
where the second expression comes by changing the variable $j$ into l=i1+k2jl=i-1+k-2j and permuting the sums. To alleviate notations, we do not specify some summation ranges as they are naturally enforced by the vanishing of the binomial and multinomial coefficients and by the requirement that their arguments are nonnegative integers. In particular, in the second expression, $k+1$ has the same parity as ili-l, which allows to reorganize the term $(\pm 1)^{k+h+1}$ as displayed.
@@ C.2.1 A binomial identity
Following [Bud22, Lemma 7], it is convenient to simplify the sum over $k$ appearing in (180). Let us first rewrite
$\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}=\binom{i-1}{l}\sum_{k}\binom{k+h}{2h+1}\binom{i-l-1}{\frac{i-l-1+k}{2}}\,$, (181)
where we recall that the sum is over nonnegative integers $k$ with the same parity as il1i-l-1. Writing j=ilj=i-l, we treat the sum over $k$ on the right-hand side by the following lemma.
@@
For any nonnegative integer $h$, there exist two polynomials $Q_{h}^{[0]}$ and $Q_{h}^{[1]}$ of degree $h+1$ such that, denoting by $\epsilon=\epsilon(j)\in\{0,1\}$ the parity of $j$, we have
$\sum_{k}\binom{k+h}{2h+1}\binom{j-1}{\frac{j-1+k}{2}}=\binom{j-1}{\frac{j-\epsilon}{2}}Q_{h}^{[\epsilon]}\left(\frac{j-\epsilon}{2}\right)$ (182)
@@
Thanks to this lemma, the sum over $k$ in (180) is expressed as
$\sum_{k}\binom{k+h}{2h+1}\binom{i-1}{\frac{i-1+k-l}{2},\frac{i-1-k-l}{2},l}=\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q_{h}^{[\epsilon]}\left(\frac{i-l-\epsilon}{2}\right)\,$, (183)
which leads to the expression
$\bar{M}_{\pm,h}=(\pm 1)^{h}\sum_{\epsilon\in\{0,1\}}(\pm 1)^{\epsilon}\sum_{i\geq 2}t_{i}\sum_{l=i+\epsilon\,\mathrm{mod}\,2}R^{\frac{i-l-1}{2}}S^{l}\binom{i-1}{\frac{i-l-\epsilon}{2},\frac{i-l+\epsilon-2}{2},l}Q^{[\epsilon]}_{h}\left(\frac{i-l-\epsilon}{2}\right)$ (184)
that we will further study in the next subsection. In passing, we note that
$Q_{h}^{[0]}(1)=0\quad\mbox{ for every }\quad h\geq 1\,$, (185)
as seen by taking $j=2$ in (182): on the left-hand side, the sum is over all positive odd $k$, and each term vanishes.
Proof of Lemma C.3.
Assume first that $j$ is even, that is, $\epsilon=0$, so that the summation is over odd $k=2k^{\prime}+1$. Then, writing $2j^{\prime}=j$, Equation (182) is rewritten as
$\sum_{k^{\prime}\geq 0}\binom{2k^{\prime}+h+1}{2h+1}\binom{2j^{\prime}-1}{j^{\prime}+k^{\prime}}=\binom{2j^{\prime}-1}{j^{\prime}}Q_{h}^{[0]}(j^{\prime})\,$, (186)
where $Q_{h}^{[0]}=Q_{h}(0,\cdot)$ is the polynomial of degree $h+1$ of [Bud22, Lemma 7] specialized at $b=0$. Alternatively, we may recover the expression of $Q_{h}^{[0]}$ by the following independent argument. First note that, in the notation of [BGM24], we have
@@
$\mathsf{A}_{\ell,k}=\frac{2k}{2\ell}\binom{2\ell}{\ell+k}\,,\quad\ell,k\in\frac{1}{2}\mathbb{Z}$ (188)
from [BGM24, Equation (3.2)], we can rewrite the left hand side of (186) in the form
$\frac{h!^{2}}{(2h+1)!}\sum_{k^{\prime}}2(k^{\prime}+1/2)p_{h}(2(k^{\prime}+1/2))\binom{2(j^{\prime}-1/2)}{j^{\prime}-1/2+k^{\prime}+1/2}\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}p_{h}(2(k^{\prime}+1/2))\,.$ (189)
@@
Taking $m=k^{\prime}+1/2$, the polynomial $p_{h}(2m)$ is of degree $h$ in $m^{2}$, so it can be expressed in the basis of the polynomials $\tilde{p}_{r}(m),0\leq r\leq h$ defined in [BGM24, Equation (2.18)], in the form $p_{h}(2m)=\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\tilde{p}_{r}(m)$, which, based on [BGM24, Equation (3.9)], yields the expression
$(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\sum_{k^{\prime}}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}\,\tilde{p}_{r}(k^{\prime}+1/2)\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime}-1/2,k^{\prime}+1/2}\,\tilde{p}_{r}(k^{\prime}+1/2)\\ =(2j^{\prime}-1)\frac{h!^{2}}{(2h+1)!}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\binom{j^{\prime}-1}{r}\binom{2j^{\prime}-2}{j^{\prime}-1}\\ =\binom{2j^{\prime}-1}{j^{\prime}}\frac{h!^{2}}{(2h+1)!}j^{\prime}\sum_{r=0}^{h}\tilde{\alpha}_{h,r}\,\binom{j^{\prime}-1}{r}$ (190)
@@
which, since $j^{\prime}=j/2$, is indeed of the wanted form $\binom{j-1}{j/2}Q_{h}^{[0]}(j/2)$ for some polynomial $Q^{[0]}_{h}$ of degree $h+1$.
@@
For $j$ odd, i.e. $\epsilon=1$, we argue in a similar (and slightly simpler) way. The summation is now over even $k=2k^{\prime}$, and we write $2j^{\prime}=j-1$, so that the left-hand side of (182) reads
$\sum_{k^{\prime}}\binom{2k^{\prime}+h}{2h+1}\binom{2j^{\prime}}{j^{\prime}+k^{\prime}}=2j^{\prime}\frac{h!^{2}}{(2h+1)!}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime},k^{\prime}}p_{h}(2k^{\prime})\,.$ (191)
@@
where we use (187) and (188) to pass to the right-hand side. This time, we take $m=k^{\prime}$ and we express the polynomial $p_{h}(2m)$ in the basis of the polynomials $p_{r}(m),0\leq r\leq h$, in the form $p_{h}(2m)=\sum_{r=0}^{h}\alpha_{h,r}p_{r}(m)$, which yields, now using [BGM24, Equation (3.7)], the expression
$2j^{\prime}\frac{h!^{2}}{(2h+1)!}\sum_{k^{\prime}}\sum_{r=0}^{h}\alpha_{h,r}\mathsf{A}_{j^{\prime},k^{\prime}}p_{r}(k^{\prime})\\ =2j^{\prime}\frac{h!^{2}}{(2h+1)!}\sum_{r=0}^{h}\alpha_{h,r}\sum_{k^{\prime}}\mathsf{A}_{j^{\prime},k^{\prime}}p_{r}(k^{\prime})\\ =2j^{\prime}\frac{h!^{2}}{(2h+1)!}\sum_{r=0}^{h}\alpha_{h,r}\binom{j^{\prime}-1}{r}\binom{2j^{\prime}-1}{j^{\prime}}\\ =\binom{2j^{\prime}}{j^{\prime}}\frac{h!^{2}}{(2h+1)!}j^{\prime}\sum_{r=0}^{h}\alpha_{h,r}\binom{j^{\prime}-1}{r}\,$, (192)
@@
which is again of the wanted form $\binom{j-1}{(j-1)/2}Q_{h}^{[1]}((j-1)/2)$ for some polynomial $Q_{h}^{[1]}$ of degree $h+1$. ∎
C.2.2 Relating the moments with the recursions determining $R,S$